On January​ 2, 2019, Kornis Corporation acquired equipment for

$300,000.

The estimated life of the equipment is 5 years or

70,000

hours. The estimated residual value is

$30,000.

What is the balance in Accumulated Depreciation on December​ 31, 2019, if Kornis Corporation uses the

double−declining−balance

method of​ depreciation?
Answer #1

Double declining rate = 200/useful life = 200/5 = 40%

Depreciation for 2019 = 300,000 * 40% = 120,000

Depreciation for 2020 = (300,000 - 120,000) * 40% = 72,000

Accumulated Depreciation = 120,000+72,000

= 192,000